the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $320,000 $520,000 $180,000 $1,020,000

From past experience, the company has learned that 30% of a month’s sales are collected in the month of sale, another 65% are collected in the month following sale, and the remaining 5% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $250,000, and March sales totaled $280,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. What is the accounts receivable balance on June 30th?
Answer #1

Answer-1)-

Statement of cash collection from debtors
Particulars April May June Total
Sales $ 320000 520000 180000 1020000
Cash Collection
February sales $ 250000*5%=12500 12500
March sales $ 280000*65%=182000 280000*5%=14000 196000
April sales $ 320000*30%=96000 320000*65%=208000 320000*5%=16000 320000
May sales $ 520000*30%=156000 520000*65%=338000 494000
June sales $ 180000*30%= 54000 54000
Total collection 290500 378000 408000 1076500

Answer-2)-

ACCOUNTS RECEIVABLE AT JUNE 30
PARTICULARS Amount
$
May sales 520000*5% 26000
June sales 180000*70% 126000
Total 152000